OpenStudy (anonymous):

The tent shown below is on display at a toy store. What is the total surface area of the tent in square feet? Assume no material is used on the bottom of the tent.

OpenStudy (anonymous):

the tent is not shown below.

OpenStudy (anonymous):

opps\

OpenStudy (anonymous):

Just take it apart, look at it in terms of a parallelogram, a rectangle, and a triangle. Solve for individual areas, add them up.

OpenStudy (anonymous):

ehhh...

OpenStudy (anonymous):

2(8 times4)?

OpenStudy (anonymous):

2(8 times 5)?

OpenStudy (anonymous):

Whole front is (not counting the triangular part on top is): 6*4 Back is the same, so: 2*6*4=48

OpenStudy (anonymous):

2(5 times 6)?

OpenStudy (anonymous):

Two sides are both 4*8, so 2*4*8=64

OpenStudy (anonymous):

2(4 times 8) :D

OpenStudy (anonymous):

The sides of the roof are just two more rectangles: 2*5*8=80

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

Finally the two triangular parts: Area of a triangle = (1/2)b*h=(1/2)*6*4=12 Both triangles: 2*12=24

OpenStudy (anonymous):

216?

OpenStudy (anonymous):

2(6 times 4) + 2(5 times 8) + 2(4 times 8) + 24?

OpenStudy (anonymous):

216?

OpenStudy (anonymous):

sum them all up, since the bottom is empty: 48+64+80+24=216

OpenStudy (anonymous):

GOOD JOB!

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I take online classes

OpenStudy (anonymous):

they seem tougher than regular classes

OpenStudy (anonymous):

I give online classes. They seem easier than regular classes.

OpenStudy (anonymous):

thats the opposite for me

OpenStudy (anonymous):

for some reason

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I got two more questions